OpenStudy (anonymous):

F=time generating force felt by the earth due to another object in space. m1=mass of the earth=5.98x10E24kg m2=mass of the in space R=distance between the earth and object (in meters) G=universal gravitational constant=6.6720x10E negative11 mE3/kg*sE2 a=radius of the earth=6400km

OpenStudy (anonymous):

F=2GM1M2a/R E3 This is the formula

OpenStudy (bahrom7893):

hey ill answer here, my formula as well as urs was wrong

OpenStudy (bahrom7893):

Okay so in the case of the moon and earth: Mm = 7.42x10^22 Me = 5.98x10^24 G = 6.6720x10^-11 a = 6400km = 6,400,000m Rm = 3.840x10^5km = 3.840x10^8m Plug them all into the formula.

OpenStudy (bahrom7893):

\[F = (2G \times M_1 \times M_2 \times a )/R^3\]

OpenStudy (bahrom7893):

F = 6.69233 x 10^34 N

OpenStudy (bahrom7893):

can u check the answer?

OpenStudy (bahrom7893):

that's for the moon part

OpenStudy (anonymous):

no my calculator not working...

OpenStudy (bahrom7893):

Oh well basically that's the formula u plug that into. Mm and Me are masses of moon and earth and are equivalent to M1 and M2

OpenStudy (anonymous):

is it Me or M1

OpenStudy (bahrom7893):

doesn't matter use whichever u want. I prefer Msube

OpenStudy (anonymous):

ok lets see if i get answer..

OpenStudy (bahrom7893):

for the sun part just replace Mm by Ms and Rm by the distance from sun to earth. convert to meters from kilometers

OpenStudy (anonymous):

is it 5.87 idk not sure..

OpenStudy (bahrom7893):

is what 5.87?

OpenStudy (anonymous):

the answer.

OpenStudy (bahrom7893):

lemme check my work

OpenStudy (bahrom7893):

hey can u start a new topic and post the following values: Mass of earth Mass of moon Mass of sun Distance b/n earth and moon Distance b/n earth and sun Universal grav. constant Radius of earth, im tired of lookin thru the posts to find values

OpenStudy (anonymous):

well im sorry, it is a lot of work i just don't know how to even do this.

OpenStudy (anonymous):

im lost

OpenStudy (bahrom7893):

the numbers u gave me earlier

OpenStudy (bahrom7893):

oh wow lol nevermind, im sick of this problem and i hate it if i cant solve something, ima do that myself.. watch out for a new topic with ur name

OpenStudy (anonymous):

sorryyy. i hate these math questions:///